OpenStudy (anonymous):

Find partial fraction of : (4s^2 + 5s + 16)/(s+7)(s^2+4)

OpenStudy (anonymous):

what a pain

OpenStudy (anonymous):

yeah man!...I solved the DFQ's with laplace transform so now I have to take the inverse laplace.

OpenStudy (anonymous):

we can do it but it is best left to a machine http://www.wolframalpha.com/input/?i=partial+fractons+%284s^2+%2B+5s+%2B+16%29%2F%28%28s%2B7%29%28s^2%2B4%29%29

OpenStudy (anonymous):

hmm..ok that confirms what I was doing because so far this is what I have: (4s^2 + 5s + 16)/(s+7)(s^2+4) =177/53(s + 7) + (K2s + K3)/s^2 + 4

OpenStudy (anonymous):

I'm trying to figure out a way to get rid of s^2 + 4 in denominator so I can solve for k2 and k3

OpenStudy (anonymous):

you are going to write \[\frac{A}{x+7}+\frac{Bx+C}{x^2+4}=\frac{4x^2+5x+16}{(x+7)(x^2+4)}\] and then \[A(x^2+4)+(Bx+C)(x+7)=4x^2+5x+16\] then multiply out and equate like coefficients

OpenStudy (anonymous):

actually we can get A right away be replacing x by -7

OpenStudy (anonymous):

the x^2 makes it impossible to get rid of x^2 + 4...yeah I did that and I got 177/53

OpenStudy (anonymous):

yup. and wolfram agrees. i wish i knew anything about these laplace transforms because they seem to come up constantly

OpenStudy (anonymous):

yeah what it all boils down to is manipulation of partial fractions usually

OpenStudy (anonymous):

but I think if we set x = 0 we can solve for C the way you set it up

OpenStudy (anonymous):

sure now that we know A it will work right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[4\times \frac{177}{53}+7C=16\] etc

OpenStudy (anonymous):

\[C=\frac{20}{53}\]?

OpenStudy (anonymous):

yep I just got that too!

OpenStudy (anonymous):

and since \[Ax^2+Bx^2=4x^2\] we can solve for B as well

OpenStudy (anonymous):

\[B=4-\frac{177}{53}\] \[B=\frac{35}{53}\]

OpenStudy (anonymous):

Cool! i get it know! Thanks

OpenStudy (anonymous):

well i can't help with the transform, but you are welcome i find this way easier than solving a 3 by 3 system of equations

OpenStudy (anonymous):

yeah the are different methods...yeah I can take it from here..thanks again!

OpenStudy (anonymous):

yw